OpenStudy (anonymous):

integral of x^2tanx/(1+cos^4x)

myininaya (myininaya):

I'm haven't' seen if it works but you can try to see if works let u=tan(x/2)

myininaya (myininaya):

that might be too ugly still

myininaya (myininaya):

mostly because that x term i think

myininaya (myininaya):

there is another weird trig substitution to try but i can't remember it

OpenStudy (rational):

looks hard for wolfram too http://www.wolframalpha.com/input/?i=%5Cint+x%5E2tanx%2F%281%2Bcos%5E4x%29

myininaya (myininaya):

maybe there is some kinda neat trick

myininaya (myininaya):

any hint @noushinrahdar given?

myininaya (myininaya):

actually I think I got it... integration by parts will be needed

myininaya (myininaya):

\[\int\limits_{}^{}x^2 \cdot \frac{\tan(x)}{1+\cos^4(x) }dx \] x^2 will be the one you are differentiating the other function you will need to integrate i preferred writing tan(x)/(1+cos^4(x)) in terms of sin and cos (of course you need to get rid of the compound fraction)

OpenStudy (anonymous):

With IBP, set \[u=x^2~~\implies~~du=2x~dx\\ dv=\frac{\tan x}{1+\cos^4x}~dx\] Integrating for \(v\): \[\begin{align*} v&=\int\frac{\tan x}{1+\cos^4x}~dx&t=\cos x\\\\ &=-\int\frac{dt}{t(1+t^4)}~dt\\\\ &=-\int\frac{A}{t}~dt-\int\frac{Bt^3+Ct^2+Dt+E}{1+t^4}~dt \end{align*}\] \[ -1=A+At^4+Bt^4+Ct^3+Dt^2+Et~~\iff~~\begin{cases}A=-1\\A+B=0\\C=D=E=0\end{cases}\] \[\begin{align*} v&=\int\frac{dt}{t}-\int\frac{t^3}{1+t^4}~dt\\\\ &=\ln|t|-\frac{1}{4}\int\frac{ds}{s}&s=1+t^4\\\\ &=\ln|t|-\frac{1}{4}\ln|s|\\\\ &=\ln|\cos x|-\frac{1}{4}\ln(1+\cos^4x) \end{align*}\] So you have \[\int\frac{x^2\tan x}{1+\cos^4x}~dx=\\ x^2\left(\ln|\cos x|-\frac{1}{4}\ln(1+\cos^4x)\right)-2\int x\left(\ln|\cos x|-\frac{1}{4}\ln(1+\cos^4x)\right)~dx\] Doesn't look elementary...

OpenStudy (anonymous):

Is this a definite integral? @noushinrahdar

myininaya (myininaya):

@SithsAndGiggles I think it is an elementary integral. I was able to integrated suggesting that way I mentioned.

myininaya (myininaya):

at least I think I did finish it let me see if i can do it again

myininaya (myininaya):

brb

OpenStudy (anonymous):

Take your time, I'm trying to see if Mathematica can spit out a result.

myininaya (myininaya):

\[\int\limits_{}^{}x^2 \cdot \frac{\tan(x)}{1+\cos^4(x)} dx\] So first I suggested writing in terms of sin and cos \[\int\limits_{}^{}x^2 \cdot \frac{\sin(x)}{\cos(x)+\cos^5(x)} dx \] Then I perform integration by parts but I took that one function to the side to find the antiderivative first of course since I didn't know it off the bat. \[\int\limits_{}^{}\frac{\sin(x)}{\cos(x)+\cos^5(x)} dx \text{ \let } u =\cos(x) => du=-\sin(x) dx \\ \int\limits_{}^{}\frac{-du}{u+u^5} du = \int\limits_{}^{} \frac{-du}{u^5(\frac{1}{u^4}+1)} \text{ \let } b=\frac{1}{u^4}+1=> \\ db=-\frac{4}{u^5} du => \frac{db}{4} =\frac{-du}{u^5} \\ \int\limits_{}^{} \frac{db}{4b}=\frac{1}{4} \ln|b|+C=\frac{1}{4} \ln|\frac{1}{u^4}+1|+C =\frac{1}{4} \ln|\frac{1}{\cos^4(x)}+1|+C\] Ok so going back to the integration by parts idea: \[\int\limits_{}^{}x^2 \frac{\sin(x)}{\cos(x)+\cos^5(x)} dx \\ =x^2 \cdot \frac{1}{4} \ln |\frac{1}{\cos^4(x)}+1| - \int\limits_{}^{}2x \cdot \frac{1}{4} \ln|\frac{1}{\cos^4(x)}+1| dx \] Now we need to perform integration by parts again. Trying to remember everything I did... So lets look at integrating that ln part \[\int\limits_{}^{} \ln |\frac{1}{\cos^4(x)}+1| dx =x \ln|\frac{1}{\cos^4(x)}+1|-\int\limits_{}^{} x \cdot 4 \frac{\sec^4(x) \tan(x)}{\frac{1}{cos^4(x)}+1}dx\] \[=x \ln|\frac{1}{\cos^4(x)}+1|-\int\limits_{}^{}4x \frac{\sin(x)}{\cos(x)+\cos^5(x)} dx \] maybe I lied

myininaya (myininaya):

hey @SithsAndGiggles can mathematica integrate ln|sec^4(x)+1| wolfram can't (I think you were right. This is definitely not an elementary integral. I think I was sleepy when I tried to find away to integrate the above. :( )

myininaya (myininaya):

http://openstudy.com/updates/512c480de4b02acc415db053 I think it is possible you were right this was maybe meant to be an definite integral

OpenStudy (anonymous):

With the absolute value, no. It just returns the input. Without, I'm getting this non-elementary expression with something called a polylogarithm and a bunch of elementary functions in terms of imaginary numbers